"An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 786 hours and a standard deviation of 80 hours. Find the probability that a random samples of 17 bulbs will have an average life between 799 and 802 hours."

1
Expert's answer
2022-01-31T18:12:32-0500

we use z calculations

z(799) = (799 - 786) / 80 = 0.1625

z(802) = (802 - 786)/ 80 = 0.2

P(799<x<802) = P(0.1625<Z<0.2)

using the normal table we have

= ( 0.57926 - 0.56356) = 0.0157 which is the required probability